# Distributed Model Checking in Graphs Classes

### v1

@@ Abstract
We show that for every first-order logic (FO) formula $\varphi$, and every graph class $\mathcal{G}$ of bounded expansion, there exists a distributed (deterministic) algorithm that, for every $n$-node graph $G\in\mathcal{G}$ of diameter $D$, decides whether $G\models\varphi$ in $\mathcal{O}(D+\log n)$ rounds under the standard CONGEST model. Graphs of bounded expansion encompass many classes of sparse graphs such as planar graphs, bounded-genus graphs, bounded-treedepth graphs, bounded-treewidth graphs, bounded-degree graphs, and graphs excluding a fixed graph $H$ as a minor or topological minor. Note that our algorithm is optimal up to a logarithmic additional term, as even a simple FO formula such as “there are two vertices of degree 3” already on trees requires $\Omega(D)$ rounds in CONGEST.
Our result extends to solving optimization problems expressed in FO (e.g., $k$-vertex cover of minimum weight), as well as to counting the number of solutions of a problem expressible in a fragment of FO (e.g., counting triangles), still running in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model. This exemplifies the vast contrast between sparse graphs and general graphs as far as CONGEST algorithms are concerned. For instance, Drucker, Kuhn, and Oshman [PODC 2014] showed that the mere problem of deciding whether a general graph contains a 4-cycle requires $\Theta(\sqrt{n}/\log n)$ rounds in CONGEST. For counting triangles, the best known algorithm of Chang, Pettie, and Zhang [SODA 2019] takes $\tilde{\mathcal{O}}(\sqrt{n})$ rounds.
Finally, our result extends to distributed certification. We show that, for every FO formula $\varphi$, and every graph class $\mathcal{G}$ of bounded expansion, there exists a certification scheme for $\varphi$ using certificates on $\mathcal{O}(\log n)$ bits. This significantly generalizes the recent result of Feuilloley, Bousquet, and Pierron [PODC 2022], which held solely for graphs of bounded treedepth.
Keywords: Sparse graphs, bounded expansion, model checking, logic, distributed computing, CONGEST model.
@@ 1 Introduction
Distributed computing in (synchronous, failure-free) networks has to cope with two obstacles, namely dilation and congestion. The dilation obstacle is that far-away nodes are required to exchange information in various network problems, e.g., constructing a Minimum Spanning Tree (MST). In this case, the diameter $D$ of the network is an inherent lower bound on the round complexity for a large class of problems. The congestion obstacle is due to limited communication bandwidth between the nodes. The standard CONGEST model [69] assumes that every communication link between adjacent nodes in the network cannot transmit more than $B$ bits in a single round of communication, where $B$ is usually set to $\Theta(\log n)$ in any $n$-node network, with the mere objective of keeping the model simple, yet informative and meaningful [^2] A bandwidth $\mathcal{O}(\log n)$ suffices to transmit elementary data items (e.g., identifiers) in a single round, and previous work has empirically demonstrated that the bounds on the throughput of CONGEST algorithms usually scale with $B$ in a straightforward manner.. The tight lower bound $\Omega(D+\sqrt{n})$ rounds in CONGEST for constructing an MST in $n$-node networks of diameter $D$ (see [70]) is a typical illustration of the price to pay for long-distance communications over links of limited bandwidth. A crucial point is that the $\Omega(D+\sqrt{n})$ rounds lower bound is not specific to MST construction but applies (up to logarithmic factors) to many problems, including approximating or verifying connectivity, spanning connected subgraph, $s$-$t$ cut, etc. (see [12]).
While the $\Omega(D+\sqrt{n})$ lower bound on the number of rounds is unavoidable for many problems even in sparse graphs [12, 18, 70], an interesting phenomenon was discovered by  [40]. It appears that, for some problems, the $\Omega(D+\sqrt{n})$ barrier could be circumvented in planar graphs. For instance, both MST and $(1+\varepsilon)$-approximate min-cut can be computed in just $\mathcal{O}(D\cdot\log^{\mathcal{O}(1)}{n})$ rounds in planar graphs by combining an efficient planar embedding algorithm [39] with the low-congestion short-cut technique [40]. The $\tilde{\mathcal{O}}(D)$ bounds [^3] The $\tilde{\mathcal{O}}$ notation, as the $\tilde{\Omega}$ and $\tilde{\Theta}$ notations, ignores factors polylogarithmic in the size $n$ of the network. on the number of rounds for MST and min-cut were extended to graphs of bounded genus [51], and graphs excluding a fixed graph as a minor [52, 41]. On the other hand, as it was observed in  [40], the bound $\tilde{\Omega}(D+\sqrt{n})$ for MST holds even when the maximum degree of a graph is at most $3$.
Besides MST and min-cut, most of the developments on the “$\tilde{\mathcal{O}}(D)$-phenomenon” focus on a very few problems like Depth-First Search [45] or approximation of maximum independent set [8], which naturally raises the question of whether this phenomenon is unique to a few optimization problems or if it extends to a broader class of problems on various classes of sparse graphs. In the centralized computing framework, a successful approach to describing general tractability results is by making use of algorithmic meta-theorems. Such theorems establish that certain families of algorithmic problems, defined by some logical conditions, can be solved “efficiently” on some classes of graphs. According to [48], algorithmic meta-theorems play an important role in the theory of algorithms as they reveal deep interplays between algorithms, logic, and combinatorics.
Algorithmic Meta-Theorems.
The most popular logics used in meta-theorems are first-order (FO) and monadic second-order (MSO) logics. By adapting known constructions (we provide a formal proof in Theorem 5), it is possible to show that deciding whether a graph of bounded degree and logarithmic diameter is not 3-colorable requires $\tilde{\Omega}(n)$ rounds. Since $3$-colorability is expressible in MSO, this immediately excludes MSO as a suitable logic for describing problems with $\tilde{\mathcal{O}}(D)$ round complexity. Nevertheless, FO remains a suitable candidate, and will be at the core of our approach.
Our understanding of model checking in classes of sparse graphs has been subject to enormous progress since the turn of the century. The first fixed-parameter tractability results (parameterized by the size of the formula) were obtained by [72] on graphs of bounded degree. [34] proved the tractability of FO model checking on planar graphs and apex-minor-free graph classes. These results were extended to graph classes with excluded minors by [27]. Progress in sparsity theory by [67, 66] have significantly extended the boundaries of FO model checking tractability on sparse graphs.
@@
Within the theory of Nešetřil and Ossona de Mendez, a crucial role is assigned to the graph classes of bounded expansion. Intuitively, a graph class $\mathcal{G}$ has bounded expansion if there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and any $r\geq 0$, every minor $H$ of $G$ obtained by contracting disjoint subgraphs of radius at most $r$ is $f(r)$-degenerate (i.e., every subgraph of $H$ has a vertex of degree at most $f(r)$). Bounded expansion can be thought of as strengthened degeneracy, which must persist even after performing “local minor operations”. For instance, the graph class containing every graph obtained from a complete graph by subdividing every edge once is sparse but not of bounded expansion. Instead, the class of planar graphs, the classes of graphs with bounded genus, the classes of graphs of bounded degree, and the classes of minor-free graphs and topologically minor-free graphs are all of bounded expansion. Interestingly, [14] demonstrated that many network models (e.g., stochastic block models) exhibit bounded expansion for some ranges of parameters, and these findings are even supported with empirical measurements on a corpus of real-world networks. All these points make graph classes of bounded expansion quite attractive from both conceptual and practical perspectives. And indeed, all the aforementioned results on model checking FO formulas can be extended graphs classes of bounded expansion, as shown by [16] who obtained an FPT algorithm for model-checking FO formulas on graph classes of bounded expansion.
The situation is not as advanced and mature in the distributed computing framework. Our work is strongly motivated by the question of whether it is possible to align the knowledge of distributed model checking with the current knowledge of centralized model checking.
Distributed Model-Checking, and Distributed Certification.
The problem of the (distributed) model checking for a property $\Pi$ on a labeled graph is formulated as follows. The algorithm checking $\Pi$ must guarantee that if the labeled graph $G$ satisfies $\Pi$, i.e., if $G\models\Pi$, then all nodes output accept, and if $G\not\models\Pi$ then at least one node must output reject. In short, a distributed algorithm $A$ decides, or checks $\Pi$ in a graph class $\mathcal{G}$ if, for every $G=(V,E)\in\mathcal{G}$,
$G\models\Pi\iff\forall v\in V,\;\mbox{$A$ accepts $G$ at node $v$}.$
As a typical example, [65] introduced the locally checkable labeling (LCL) problems as a formalization of problems for which the global correctness of a solution is determined as the conjunction of local criteria (e.g., maximal independent set, or proper coloring). LCL plays a central role in distributed network computing [69]. In bounded degree graphs, an LCL can be checked in a constant number of rounds in CONGEST. However, simple FO properties such as $C_{4}$-freeness require $\tilde{\Omega}(\sqrt{n})$ rounds to be checked in general graphs in CONGEST [15].
As a byproduct of our techniques developed for FO model checking in CONGEST model, we are also questioning the existence of meta-theorems susceptible to extend the known results about distributed certification [30, 46, 57], which plays a major role in the design of fault-tolerant algorithms. Distributed certification can be conceptually thought of as a non-deterministic variant of distributed model checking. Roughly, a certification scheme for a graph property $\Pi$ is a pair prover-verifier. The prover is a computationally unlimited oracle that assigns a certificate $c(v)$ to each node $v$. The verifier is a distributed algorithm whose role is to check whether the certificates assigned by the prover establish evidence that the network satisfies $\Pi$. The scheme certifying $\Pi$ must guarantee that if $G\models\Pi$ then the prover can assign certificates such that the verifier accepts at each and every node, and if $G\not\models\Pi$ then, for every certificate assignment by the prover, the verifier rejects in at least one node (i.e., the verifier cannot be fooled by the prover on illegal instances). In short, a scheme certify $\Pi$ in a graph class $\mathcal{G}$ if, for every $G=(V,E)\in\mathcal{G}$,
$G\models\Pi\iff\exists c:V\to\{0,1\}^{\star},\;\forall v\in V,\;\mbox{verifieraccepts $(G,c)$ at node $v$}.$
Distributed certification actually provides a concrete mechanism enabling the nodes of a network to quickly determine whether the system entered an illegal state, and to fix it (a node at which the verifier rejects can launch a recovery procedure). Indeed, for very many problems (e.g., MST construction), the certificates can be collectively computed by the nodes themselves, in parallel to computing a solution, with negligible extra cost.
In distributed certification, it is desirable that the verifier perform as quickly as possible. The focus is actually on verifiers such that the decision (accept or reject) taken at every node $v$ depends only on the certificates assigned by the prover to the nodes in the closed neighborhood $N[v]$ of node $v$, that is, on verifiers performing in a single round of communication. The main complexity measure for a certification scheme is the size $|c(v)|$ of the certificate assigned by the prover to every node $v$ of legal instances. It was proved that every FO property could be distributedly certified using certificates of $\mathcal{O}(\log n)$ bits on graphs with bounded treedepth [22], and that every MSO property can be certified using certificates of $\mathcal{O}(\log^{2}n)$ bits on graphs with bounded treewidth [32] and cliquewidth [31]. Note that $\mathcal{O}(\log n)$-bit certificates enable the verifier to perform a single round of communication in CONGEST for exchanging the certificates between neighbors, whereas $\mathcal{O}(\log^{2}n)$-bit certificates require $\Theta(\log n)$ rounds for exchanging the certificates between neighbors.
@@ 1.1 Our Results
1.1.1 Distributed Model Checking
Our first main result states that every graph property expressible as an FO formula can be decided in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model in $n$-node graphs of bounded expansion, and diameter $D$. This result actually holds for labeled graphs as well. A labeled graph is a pair $(G,L)$ where $G=(V,E)$ and $L:V\to 2^{\Lambda}$ assigns to each node a set of labels taken from a finite set $\Lambda$. FO formulas on graphs can be extended to FO formulas on labeled graphs by merely introducing a unary predicate $\textsf{lab}_{\lambda}(x)$ for every $\lambda\in\Lambda$, which holds if and only if $\lambda$ is one of the labels of vertex $x$ in $L$.
Theorem 1.
@@ 1.1.1 Distributed Model Checking
Let $\Lambda$ be a finite set. For every FO formula $\varphi$ without free variables on $\Lambda$-labeled graphs, and for every class of $\Lambda$-labeled graphs $\mathcal{G}$ of bounded expansion, there exists a distributed algorithm which, for every $n$-node network $G=(V,E)\in\mathcal{G}$ of diameter $D$, and every $L:V\to 2^{\Lambda}$, decides whether $(G,L)\models\varphi$ in $\mathcal{O}(D+\log n)$ rounds under the CONGEST model.
The constant hidden in the big-O notation depends on the length $|\varphi|$ of the formula $\varphi$, on the set $\Lambda$, and on the function controlling the expansion of $\mathcal{G}$, but not on the graph $G$ nor on the labeling $L$. Theorem 1 complements the previous work on MST and $s,t$-cut of [39, 40, 41, 52] circumventing the $\tilde{\Omega}(\sqrt{n})$ barrier on the congestion of distributed algorithms in CONGEST in two directions.
First, it holds for the large class of problems definable by FO formulas. By making use of the expressive power of FO, Theorem 1 implies that the bound $\mathcal{O}(D+\log n)$ on the number of rounds under the CONGEST model on graphs of bounded expansion holds for various problems, such as deciding whether a graph contains (or excludes) a cycle of a fixed length, or, more generally, whether it contains a subgraph isomorphic to a fixed graph of constant size, a connected dominating set of a fixed size, or a Steiner tree of a fixed size.
Second, Theorem 1 holds for a significantly more general class of sparse graphs than minor-free graph. Moreover, as we already mentioned, the lower bound $\tilde{\Omega}(D+\sqrt{n})$ holds for MST (which is not FO-definable) when the maximum degree of a graph is at most $3$. On the other hand, graphs of bounded degree, as well as a more general class of graphs excluding some fixed graph as a topological minor, are of bounded expansion.
The dependence in $D$ in Theorem 1 cannot be avoided, as witnessed by the mere FO property “there are at least two nodes of degree 3” in binary trees. The dependence on the size $n$ of the networks is however drastically reduced from $\Omega(\sqrt{n})$ rounds in general graphs (e.g., for the FO formula $C_{4}$-freeness [15]) to just $\mathcal{O}(\log n)$ rounds in graphs of bounded expansion. Note also that bounded degeneracy would not suffice to establish our result, as we show that the mere FO formula $C_{6}$-freeness requires $\tilde{\Omega}(\sqrt{n})$ rounds to be checked in CONGEST, in graphs of degeneracy 2 and of bounded diameter (see 7). Moreover, our result is the best that one can hope for regarding the class of problems, in the sense that it cannot be extended to, e.g., MSO. Indeed, we show that non-3-colorability requires $\tilde{\Omega}(n)$ rounds to be checked in bounded degree graphs with logarithmic diameter (see Theorem 5).
1.1.2 Counting and Optimization
Our result extends beyond model checking, in two directions. Firstly, we consider the problem of counting solutions of FO formulas with no free variables. This includes triangle counting, and more generally counting copies of some given subgraph.
Theorem 2.

### v2

# Pedro Montealegre Universidad Adolfo Ibañez, Santiago, Chile Ivan Rapaport Universidad de Chile, Santiago, Chile Ioan Todinca LIFO, Université d’Orléans and INSA Centre-Val de Loire, Orléans, France

@@ Abstract
The sparsity theory of Nešetřil and Ossona de Mendez provides generic combinatorial and algorithmic methods suitable for many classes of sparse graphs. As such, it has already driven major advances in model checking, parameterized complexity, and approximation algorithms. We show that this theory is equally well suited to the design of distributed algorithms, providing the first generic algorithmic results for distributed first-order (FO) model checking on graphs of bounded expansion. The latter form a rich graph family which subsumes planar, bounded-genus, bounded-treedepth, bounded-treewidth, and bounded-degree graphs, as well as graphs excluding a fixed minor or topological minor, sparse Erdős–Rényi graphs (a.a.s.), and several network models such as stochastic block models for suitable parameter ranges.
Our main results are the following algorithmic meta-theorems for distributed algorithms in the standard CONGEST model on every graph class $\mathcal{G}$ of bounded expansion.
  • First, we resolve an open problem originally posed by Nešetřil and Ossona de Mendez (Distributed Computing 2016), and reiterated by Pilipczuk, Siebertz, and Toruńczyk (LICS 2018). Recall that a formula $\varphi(x)$ is local if the satisfaction of $\varphi(x)$ depends only on the $r$-neighbourhood of its free variable $x$, for some fixed $r$. For instance, the formula “$x$ belongs to a triangle” is local. We show that, for every local FO formula $\varphi(x)$, there exists a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$, identifies all vertices $v\in V(G)$ such that $G\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds.
  • Second, we show that, for every FO formula $\varphi$, there is a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$, decides whether $G\models\varphi$, in $\mathcal{O}(D+\log n)$ rounds, where $D$ is the diameter of $G$. Note that the dependence on $D$ is unavoidable, as even the simple FO formula “there exist two vertices of degree 3” needs $\Omega(D)$ rounds in CONGEST, even on trees.
The techniques behind these meta-theorems extend to distributed counting, optimization, and certification problems on graphs of bounded expansion.
@@ 1 Introduction
1.1 Context and Objective
What is a sparse graph? The answer to this question depends on the perspective. For instance, the graph class in which every graph is obtained from a complete graph by subdividing every edge once is “combinatorially sparse” (it has a number of edges linear in the number of vertices), but is not necessarily “effectively sparse” from an algorithmic perspective. The sparsity theory by Nešetřil and Ossona de Mendez [53, 54] provided us with an approach of sparsity that can be effectively used algorithmically.
1.1.1 Graphs of Bounded Expansion
@@
Within the sparsity theory, a crucial role is assigned to the graph classes of bounded expansion. Intuitively, a graph class $\mathcal{G}$ has bounded expansion if there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and any $r\geq 0$, every minor $H$ of $G$ obtained by contracting disjoint subgraphs of radius at most $r$ is $f(r)$-degenerate (i.e., every subgraph of $H$ has a vertex of degree at most $f(r)$). Bounded expansion can be thought of as strengthened degeneracy, which must persist even after performing “local minor operations”. For instance, the aforementioned graph class containing every graph obtained from a complete graph by subdividing every edge once is not of bounded expansion.
Graph classes with bounded expansion form a rich graph family (see Figure 1). For instance, the class of planar graphs, the class of graphs with bounded genus, the class of graphs of bounded degree, and the class of minor-free graphs and topologically minor-free graphs are all of bounded expansion, as well as sparse Erdős–Rényi graphs (a.a.s.). Moreover, [15] demonstrated that many network models (e.g., stochastic block models) exhibit bounded expansion for some ranges of parameters, and these findings are even supported with empirical measurements on a corpus of real-world networks. All these points make graph classes of bounded expansion quite attractive from both conceptual and practical perspectives. In particular, it has already driven major advances in model checking [18], parameterized complexity [53], kernelization [16, 20] and approximation algorithms [43].
Bounded degeneracyBounded expansionExcluded topological subgraphExcluded minorBounded degreeBounded genusBounded treewidthPlanarBounded treedepth
Figure 1: Different classes of sparse graphs, with their mutual inclusion relations. A arrow from a class $\mathcal{G}$ to a class $\mathcal{G}^{\prime}$ indicates that $\mathcal{G}^{\prime}$ is (strictly) included in $\mathcal{G}$.
1.1.2 Distributed Computing in Graphs of Bounded Expansion
Nešetřil and Ossona de Mendez also showed that bounded expansion can also be approached from a distributed computing perspective (see [55]). Their finding is based on the following characterization [54]. A class $\mathcal{G}$ of graphs has bounded expansion if and only if there exists a function $g:\mathbb{N}\to\mathbb{N}$ such that, for every $G\in\mathcal{G}$, and for every $k\in\mathbb{N}$, there exists a coloring of the vertices of $G$ with colors in $[g(k)]=\{1,\dots,g(k)\}$ such that, for every set $U\subseteq[g(k)]$ of cardinality at most $k$, the subgraph of $G$ induced by the nodes colored with the colors in $U$ has treedepth at most $|U|$. (Recall that a graph $G=(V,E)$ has treedepth at most $d$ if there exists a forest $F=(V,E^{\prime})$ of rooted trees, on the same set of vertices as $G$, such that (1) every tree in $F$ has depth at most $d$, and (2) for every edge in $E$, one of its end points is an ancestor of the other in some tree of $F$.)
Nešetřil and Ossona de Mendez [55] established that for every graph class $\mathcal{G}$ of bounded expansion $f$, there exists a distributed algorithm which, for every positive integer $k$, and for every $G\in\mathcal{G}$, computes a $g(k)$-coloring of the nodes of $G$ satisfying the desired treedepth property, in $\mathcal{O}(\log n)$ rounds in the CONGEST model [^2] The CONGEST model [57] is a standard distributed computing model in which the nodes of the graph perform in a sequence of synchronous rounds, during each of which every pair of adjacent nodes can exchange a message on $\mathcal{O}(\log n)$ bits, e.g., a constant number of node identifiers.. The function $g$ depends on the given bound $f$ on the expansion of $\mathcal{G}$.
1.1.3 Deciding Local First-Order Formulas in CONGEST
The fact that a coloring satisfying the requirement of bounded expansion can be computed in a distributed manner yields the question of whether this ability can be used as a building block for solving “hard problems” efficiently in CONGEST, in graph classes of bounded expansion. In the context of CONGEST, “hard problems” are usually thought as problems requiring at least $\mbox{poly}(n)$ rounds to be solved, whereas “easy problems” are problems solvable in $\mbox{poly}\log(n)$ rounds. As a concrete illustration, it is known [17] that deciding $C_{4}$-freeness is hard in CONGEST, as it requires $\tilde{\Omega}(\sqrt{n})$ rounds — in the distributed setting, deciding a property asks all nodes to accept if the property is satisfied, and at least one node to reject if the property is violated. Is it possible to overcome the $\tilde{\Omega}(\sqrt{n})$ lower bound for deciding $C_{4}$-freeness in graphs of bounded expansion?
Nešetřil and Ossona de Mendez [55] actually answered positively to this question, by highlighting a general approach for deciding $H$-freeness, in $\mathcal{O}(\log n)$ rounds, for any fixed (connected) graph $H$. In fact, this approach can even be used to mark all nodes $v$ such that $v$ belongs to some subgraph of $G$ isomorphic to $H$. However, they asked a broader and more challenging question, reiterated by Pilipczuk, Siebertz, and Toruńczyk [59]. Recall that a formula $\varphi(x)$ is local if the satisfaction of $\varphi(x)$ depends only on the $r$-neighborhood of its free variable $x$, for some fixed $r$. For instance, in graphs, the formula “node $x$ belongs to a 4-cycle” is indeed local. The question in [55, 59] goes way beyond deciding $H$-freeness. It can be formulated as follows:
Open Problem (Nešetřil and Ossona de Mendez [55]):
For every local first-order (FO) formula $\varphi(x)$, and for every graph class $\mathcal{G}$ of bounded expansion, is there a distributed algorithm that, for every $n$-node graph $G\in\mathcal{G}$, marks all vertices $v\in V(G)$ such that $G\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds in the CONGEST model?
An example of such local formula $\varphi(x)$ is, for fixed positive integers $d$ and $k$, marking all nodes $v$ such that the distance-$d$ neighborhood of $v$ in $G$ has a dominating set of size at most $k$.
@@ 1.2 Our Results
Our first main result is a positive answer to the open problem in [55, 59]. That is, we show that, for every local FO formula $\varphi(x)$, and for every graph class $\mathcal{G}$ of bounded expansion, there exists a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$, marks all vertices $v\in V(G)$ such that $G\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds in the CONGEST model. In fact, this result also holds for FO formulas on labeled graphs, i.e., graphs in which every node is provided with labels drawn from a finite set $\Lambda$. That is, the formulas may also include unary predicates on node labels. Let us call a $\Lambda$-labeled graph any graph whose nodes are label by labels in $\Lambda$. A $\Lambda$-labeled graph is thus a pair $(G,\ell)$ where $G=(V,E)$ is a graph, and $\ell:V\to 2^{\Lambda}$ is a label-assignment to the nodes of $G$. Our first main theorem is the following:
Theorem 1.1.
@@
Let $\Lambda$ be a finite set. For every local FO formula $\varphi(x)$ on $\Lambda$-labeled graphs, and for every class of graphs $\mathcal{G}$ of bounded expansion, there exists a deterministic algorithm that, for every $n$-node network $G=(V,E)\in\mathcal{G}$, and every $\ell:V\to 2^{\Lambda}$, marks all vertices $v\in V$ such that $(G,\ell)\models\varphi(v)$, in $\mathcal{O}(\log n)$ rounds under the CONGEST model.
Before summarizing our techniques used to solve the open problem in [55, 59], it is worth mentioning that we extend this result in several directions.
1.2.1 Extension 1: Deciding FO formulas
An immediate corollary of our result on local FO formulas is that, for every formula $\varphi$ such as $\varphi=\forall x\;\psi(x)$ where $\psi(x)$ is a local FO formula (e.g., $H$-freeness), and for every graph class $\mathcal{G}$ of bounded expansion, deciding whether $G\models\varphi$ can be done in $\mathcal{O}(\log n)$ rounds in the CONGEST model (by merely applying our algorithm to $\neg\psi(x)$, and letting every marked node reject). However, Gaifman’s theorem [29] tells us that not all FO formulas are as simple as that. For instance, even the simple FO formula “there exist two vertices of degree 3” needs $\Omega(D)$ rounds to be checked in CONGEST, where $D$ denotes the diameter of the graph — this lower bound holds even for trees. In other words, a dependence on $D$ is unavoidable, even for trees. Nevertheless, our second main result states that, for every FO formula $\varphi$, and for every graph class $\mathcal{G}$ of bounded expansion, there exists a deterministic algorithm that, for every $n$-vertex graph $G\in\mathcal{G}$ of diameter $D$, decides whether $G\models\varphi$, in $\mathcal{O}(D+\log n)$ rounds. Again, our result holds for labeled graphs too, and is formally stated as follows:
Theorem 1.2.